You flip a coin 10 times.What is the P(8H)?What is the P(>8H)?
Please answer this question.

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

You flip a coin 10 times.What is the P(8H)?
P(8heads in 10) = 10C8(1/2)^8(1/2)^2 = 45/2^10 = 0.0439
---------------------
What is the P(>8H)?
P(9<= x <=10) = (10C9 + 10C10)/2^10 = (10+1)/2^10 = 11/2^10 = 0.0107
======================================
Cheers,
Stan H.

Answer by solver91311(24713)



The probability of successes out of trials where is the probability of success on any given trial is given by:



Where is the number of ways to select things out of things and is equal to .

So for the first part of your question, you need to calculate:



Because you want the probability of 8 heads in 10 flips where the probability of getting heads is

The probability of more than successes out of trials where is the probability of success on any given trial is given by:



So, you need to calculate:



Use MS Excel to check your work. Pick any cell on a blank worksheet and type in:



for the first part and



for the second part.


John